i need help solving this equation for y: 5{-7[6y+9(7-y)-5]+y} = 6(y + 9)
i tried doing the opposite of order of operations. i've tried distubuting everything. if u could help me, i would appreciate it a lot. thank you. :) -Guadalupe

Answer by stanbon(75887)

5{-7[6y+9(7-y)-5]+y} = 6(y + 9)
5[-7[6y+63-9y-5]+y = 6y+54
5[-7[-3y+58]+y = 6y+54
5[21y-406]+y = 6y+54
105y-2030+y = 6y+54
106y-2030 = 6y+54
100y=2084
y=20.84 or 521/25
Cheers,
Stan H.